Find the volume of the ice cream cone that is bounded above by the half sphere z = sqrt( 4− x^2 − y^2 ) + 4 and the cone z= 2sqrt( x^2 + y^2 ). Transformation to cylindrical coordinates could be easier.
Answer #1

n .. a(x,y) = V4 - x2 - y2 . b(x,y) = 2 x2 + y2 . 4 fizə 5 + Input...AZ z = 3 اڑ ج ylk of X 21 5, is the part of the upper hemisphere 4-nr-yor S in the cone 2/arty? Intersection S and P2 in theCite transform the system to cylndbrical coordinate system by the tram formation (9,82) = (1,0,2) y = r sina Z N= acona Hy Z2. bu () ()? qidade do 1671 1555 (4-22 r dy drdo And Q=0 2=0 2= is 2/15 27 2/5 Hr dr T5 * 20/214- 4r dr da 1) ruar VE o Jo 6